OpenStudy (anonymous):

1.What is the slope for the following table of values? Gallons of has used: 1,2,3,4 Miles driven: 26,52,78,104 2. Write an equation in y= mx + b form for the previous problems table. A: y=26x B: y=26x + 26 C: y=26x + 1 D: y=26x - 1

OpenStudy (anonymous):

Take two pairs of points, (1,26) and (2,52). slope = change in y=values/change in x-values = [(52-26)/(2-1)] = 26/1 = 26 The slope of the line is 2. The equation of the line is y = mx + m, with m, the slope = 26 so y = 26x + b Let x = 1 and y = 26, so 26 = 26(1) + b, b = 0 Equastion of the line is y = 26x.